# Partial Information Rate Decomposition

### v1

@@ Abstract
Figure 1: Decomposition of the multivariate information shared instantaneously and dynamically in a network Gaussian system (a) with instantaneous (red) and time-lagged (internal: orange; causal: green) simulated interactions. In the absence of instantaneous effects, the zero-lag multivariate information is modulated non-trivially by time-lagged effects and is predominantly redundant (b, left), while the dynamic information reflects the strength of time-lagged effects and is predominantly synergistic (b, right). When a transition from pure instantaneous to pure time-lagged effects is simulated (c), PID (solid lines) fails in detecting the trend and nature of multivariate interactions, while PIRD reveals the expected emergence of dynamic coupling (dashed, black) and net synergy (dashed, blue vs. red).
@@
As an application to real data, we consider an exemplary case study in climate science, i.e. the network of interactions among the most representative indices descriptive of El Niño and the Southern Oscillation (ENSO). ENSO is a periodic fluctuation in the sea surface temperature and air pressure of the atmosphere overlying the equatorial Pacific Ocean, which is considered as the most prominent interannual climate variability on Earth McPhaden et al. (2006). Since the exact initiating causes of an ENSO warm or cool events are not fully understood, it is important to analyze the statistical relation between its two main components, i.e., atmospheric pressure and sea surface temperature. Such components are measured respectively by NINO34 (the East Central Tropical Pacific sea surface temperature anomaly, also called El Niño) and SOI (Southern Oscillation Index, the standardized difference in surface air pressure between Tahiti and Darwin), and are dynamically related to several other indexes that represent large scale climate patterns Chang et al. (2003); Silini et al. (2023). The resulting network of causal interactions is densely connected through several feedback loops of inter-annual variations Silini et al. (2023), suggesting the importance of studying high-order interactions with a comprehensive dynamic approach. Here, in addition to NINO34 and SOI we considered three more indices which can have a remarkable impact on ENSO as a result of high-order effects, i.e. TSA (Tropical Southern Atlantic Index), PDO (Pacific Decadal Oscillation), and NTA (North Tropical Atlantic) Stramaglia et al. (2024).
The analyzed climate indices are taken from a public database Silini et al. (2023), of which we consider the series SOI, NINO34, TSA, PDO and NTA measured with a monthly sampling rate during the period 1950-2016 (792 data points) for which all time series values are available. The series were detrended and deseasonalized, and then analyzed under the assumption of Gaussianity so as to identify the terms relevant to the PID and PIRD formulations (the Gaussian assumption was adopted also in in Silini et al. (2023); Stramaglia et al. (2024)). The series were analyzed in triplets, considering SOI as the target and all possible pairs of other series as sources; for each triplet, PIRD and zero-lag PID were computed respectively from the parameters of a vector autoregressive (VAR) model Lütkepohl (2005) (least squares estimation, model order set according to the Akaike information criterion Faes et al. (2012)) and a linear regression model identified on the three series (least squares estimation). To assess the significance of temporal correlations, the PIRD was computed both on the original time series and on surrogate series for which the static MI was preserved (random shuffling of the samples of each series, with the same random permutation applied to all series). We find that the pairs of sources including NINO34 share the highest amounts of information with the target SOI (Fig. 2a), and that such information arises from a unique contribution of NINO34 (Fig. 2b); unique effects are noticeable also when considering NTA ($X_{2}$ in the triplets 3,5,6, Fig. 2b). The index NTA contributes significantly also to high-order interactions (both redundant and synergistic, triplets 3,5,6 in Fig. 2c). Crucially, the importance of NTA is elicited only using PIRD and, more generally, surrogate data demonstrates that temporal correlations play a fundamental role in information decomposition and cannnot be disregarded. These results confirm the main role played by El Niño temperature anomalies on athmospheric pressure and ENSO dynamics Silini et al. (2023); Stramaglia et al. (2024), and documents the presence of dynamic many-body effects involving the north tropical atlantic climate mode and the ENSO components.
@@
In summary, the proposed framework offers a novel comprehensive method to parcel out the information shared by dynamic network systems into amounts quantifying unique and high-order effects. We posit that this framework generalizes and unifies previous PID theories applied to multivariate time series, establishing a new approach for the analysis of complex networks that reconciles information-theoretic and dynamical systems perspectives. Indeed, as it explicitly takes the full temporal statistical structure of random processes into account, PIRD overcomes the intrinsic limitation of PID of being unable to deal with temporally correlated variables. Moreover, the proposed spectral formulation of PIRD based on redundancy rates defined in the frequency domain allows to decompose dynamic information considering predetermined oscillations with specific meaning for the analyzed dynamic network. We expect that these properties, together with the algorithmic reliability and low computational demand of the implementation proposed here for PIRD, will open new avenues for the analysis of real-world networks with dynamic oscillatory behavior.
Data and code availability
The code to simulate and analyze data is available at www.lucafaes.net/PIRD.html. Climate data as described in Silini et al. (2023) can be downloaded at the NOAA website https://psl.noaa.gov/data/climateindices/list/.

### v3

@@ Abstract
Figure 1: Decomposition of the multivariate information shared instantaneously and dynamically in a network Gaussian system with instantaneous (red) and time-lagged (internal: orange; causal: green) simulated interactions (a). In the absence of instantaneous effects, the zero-lag multivariate information is modulated non-trivially by time-lagged effects and is predominantly redundant (b, left), while the dynamic information reflects the strength of time-lagged effects and is predominantly synergistic (b, right). When a transition from pure instantaneous to pure time-lagged effects is simulated (c), PID (solid lines) fails in detecting the trend and nature of multivariate interactions, while PIRD reveals the expected emergence of dynamic coupling (dashed, black) and net synergy (dashed, blue vs. red).
The system is analyzed by computing both PID and PIRD from the true values of the parameters imposed in (8); details of the computation are reported in Sparacino et al. (2025a), while the code to simulate and analyze data is available at www.lucafaes.net/PIRD.html. The analysis is performed as a function of a modulation parameter $d\in[0,1]$ that acts on the strength of instantaneous and time-lagged effects. First, we study (8) in the absence of instantaneous effects ($R_{U_{1}U_{2}}=R_{WU_{k}}=0$), letting the strength of causal interactions to increase progressively ($c_{1}=c_{2}=d$), and the strength of internal dynamics to decrease progressively ($a_{1}=a_{2}=0.8(1-d)$), with fixed coupling between the sources ($b_{1}=b_{2}=0.1$). The analysis (Fig. 1b) documents that the target shares information at lag zero with the sources as soon as time-lagged effects are set ($I(Y(t_{n});X(t_{n}))>0$ when $d>0$) even if instantaneous effects are kept to zero; this information does not exhibit a monotonic trend, and PID incorrectly interprets it as prevalently redundant ($R(Y(t_{n});X(t_{n}))>S(Y(t_{n});X(t_{n}))$). On the other hand, PIRD captures the increasing rate of information shared by the target and source processes, and documents the expected emergence of net synergy consequent to the increasingly prevailing strength of the common target effect of $X_{1}$ and $X_{2}$ on $Y$; the unique terms are null in this simulation where the two sources contribute equally to the target. In the second simulation setting, the system is investigated reproducing a transition from purely instantaneous to purely time-lagged interactions among the processes, achieved decreasing progressively the strength of zero-lag effects ($R_{U_{1}U_{2}}=R_{WU_{2}}=0.25(1-d)$, $R_{WU_{1}}=0.5(1-d)$) while increasing the strength of internal dynamics and causal interactions ($a_{1}=a_{2}=0.2d$, $b_{1}=b_{2}=0.1d$, $c_{1}=c_{2}=0.6d$). The results in Fig. 1c confirm that the PIRD reduces to the PID in the absence of time-lagged effects ($d=0$), but the injection of temporal correlations into the system makes the two approaches to differ substantially. The information shared between $Y(t_{n})$ and $\{X_{1}(t_{n}),X_{2}(t_{n})\}$ decreases progressively together with its unique, redundant and synergistic contributions, preventing meaningful interpretations. On the other hand, PIRD allows describing expected trends, such as the initial decline of the overall information rate between $Y$ and $\{X_{1},X_{2}\}$ due to the decreasing unique contribution of $X_{1}$, and its subsequent increase due to the rise of both dynamic redundancy and synergy; net synergy emerges for $d>0.7$ when common target structures prevail over common drive ones.
@@
As an application to real data, we consider an exemplary case study in climate science, i.e. the network of interactions among the most representative indices descriptive of El Niño and the Southern Oscillation (ENSO), a periodic fluctuation in the sea surface temperature and air pressure of the atmosphere overlying the equatorial Pacific Ocean constituting the most prominent interannual climate variability on Earth McPhaden et al. (2006). Since the exact initiating causes of an ENSO warm or cool events are not fully understood, it is important to analyze the statistical relation between its two main components, i.e. the East Central Tropical Pacific sea surface temperature anomaly (known as El Niño, NINO34) and the standardized difference in surface air pressure between Tahiti and Darwin (Southern Oscillation Index, SOI). These processes are dynamically related to several other indexes that represent large scale climate patterns Chang et al. (2003); Silini et al. (2023), forming a network of causal interactions densely connected through several feedback loops of inter-annual variations Silini et al. (2023). Here, in addition to NINO34 and SOI we considered three more climate indices which can have a remarkable impact on ENSO as a result of high-order effects, i.e. TSA (Tropical Southern Atlantic Index), PDO (Pacific Decadal Oscillation), and NTA (North Tropical Atlantic) Stramaglia et al. (2024). All indices are available at the NOAA website (https://psl.noaa.gov/data/climateindices/list/) in the form of time series measured with a monthly sampling rate during the period 1950-2016 (792 data points). The series were detrended and deseasonalized, and then analyzed under the assumption of Gaussianity so as to identify the terms relevant to the PID and PIRD formulations (the Gaussian assumption was adopted also in in Silini et al. (2023); Stramaglia et al. (2024)). The series were analyzed in triplets, considering SOI as the target and all possible pairs of other series as sources; for each triplet, PIRD and zero-lag PID were computed respectively from the parameters of a vector autoregressive (VAR) model Lütkepohl (2005) (least squares estimation, model order set according to the Akaike information criterion Faes et al. (2012)) and through simple linear regression (least squares estimation). To assess the significance of temporal correlations, the PIRD was computed both on the original time series and on surrogate series for which the static MI was preserved (random shuffling of the samples of each series, with the same random permutation applied to all series).
@@
We found that the pairs of sources including NINO34 share the highest amounts of information with the target SOI (Fig. 2a), and that such information arises from a unique contribution of NINO34 (Fig. 2b); unique effects are noticeable also when considering NTA ($X_{2}$ in the triplets 3,5,6, Fig. 2b). The index NTA contributes significantly also to high-order interactions (both redundant and synergistic, triplets 3,5,6 in Fig. 2c). Crucially, the importance of NTA is elicited only using PIRD, as synergistic PID values are negligible. More generally, surrogate data demonstrate that temporal correlations play a fundamental role in information decomposition and cannot be disregarded; the finding that PIRD measures can be both larger and smaller than the corresponding PID measures documents that time-lagged effects mix with each other and with instantaneous effects in a non-trivial way. These results confirm the main role played by El Niño temperature anomalies on athmospheric pressure and ENSO dynamics Silini et al. (2023); Stramaglia et al. (2024), and suggest the importance of studying high-order interactions with a dynamic approach in this application. In fact, the redundant and synergistic dynamic effects emphasized only by PIRD among the two ENSO components and the NTA reveal the existence of wide climatological patterns which extend on a global scale involving the west southern pacific, east central pacific and north tropical atlantic climate modes.
In summary, the proposed framework offers a novel comprehensive method to parcel out the information shared by dynamic network systems into amounts quantifying unique and high-order effects. We posit that this framework generalizes and unifies previous PID theories applied to multivariate time series, establishing a new approach for the analysis of complex networks that reconciles information-theoretic and dynamical systems perspectives. Indeed, as it explicitly takes the full temporal statistical structure of random processes into account, PIRD overcomes the intrinsic limitation of PID of being unable to deal with temporally correlated variables. The proposed spectral formulation of PIRD is valid for Gaussian systems and rests on the degree to which the system at hand can be described exhaustively by spectral analysis or linear models. This restriction does not invalidate the analysis of nonlinear systems, as stationary random processes are guaranteed to have a linear (albeit infinite-order) representation even if the generating mechanism of the process is nonlinear Hannan (1979). Moreover, the formulation based on redundancy rates defined in the frequency domain allows to decompose dynamic information considering predetermined oscillations with specific meaning for the analyzed dynamic network Sparacino et al. (2025a). We expect that these properties, together with the algorithmic reliability and low computational demand of the implementation proposed here for PIRD, will open new avenues for the analysis of real-world networks with dynamic oscillatory behavior.